# Dragging with Geometry: From Pixels to Geometry-Guided Image Editing

### v1

@@ Abstract
Interactive point-based image editing serves as a controllable editor, enabling precise and flexible manipulation of image content. However, most drag-based methods operate primarily on the 2D pixel plane with limited use of 3D cues. As a result, they often produce imprecise and inconsistent edits, particularly in geometry-intensive scenarios such as rotations and perspective transformations. To address these limitations, we propose a novel geometry-guided drag-based image editing method—GeoDrag, which addresses three key challenges: 1) incorporating 3D geometric cues into pixel-level editing, 2) mitigating discontinuities caused by geometry-only guidance, and 3) resolving conflicts arising from multi-point dragging. Built upon a unified displacement field that jointly encodes 3D geometry and 2D spatial priors, GeoDrag enables coherent, high-fidelity, and structure-consistent editing in a single forward pass. In addition, a conflict-free partitioning strategy is introduced to isolate editing regions, effectively preventing interference and ensuring consistency. Extensive experiments across various editing scenarios validate the effectiveness of our method, showing superior precision, structural consistency, and reliable multi-point editability. The code will be available on https://github.com/xinyu-pu/GeoDrag.
@@ 1 Introduction
Image editing [1, 2, 3] has seen remarkable progress in recent years, largely driven by the emergence of powerful generative models [4, 5, 2, 6, 7]. Among these, text-guided image editing [8, 9, 7, 10] has become a widely adopted paradigm, allowing users to modify images using natural language prompts. While expressive and flexible, this approach often falls short in providing fine-grained spatial control—especially when precise, localized, or geometry-sensitive edits are required [11, 3, 12, 13, 14, 15, 16].
@@
To overcome these limitations, point-based image editing [11, 3, 12, 13, 14, 15, 16] has emerged as a powerful alternative, gaining traction for its user-friendly, precise manipulation. By enabling users to specify handle-to-target point pairs, methods such as DragGAN [11] and its successors [3, 12, 13, 14, 15, 16, 17] offer intuitive and precise image manipulation. They often adopt a two-step pipeline: (1) motion supervision, dragging the handle point toward the target, and (2) point tracking, monitoring the handle’s updated position during editing.
@@
Motivation. Despite their success, existing point-based methods often rely on iterative gradient-based optimization, which is computationally intensive and impractical for real-time use. To improve efficiency, recent approaches like FastDrag [18] and RegionDrag [19] propose one-step editing via latent relocation based on dense displacement fields constructed over user-specified regions. However, these methods operate purely in the 2D pixel plane, and ignore the underlying 3D scene geometry. This becomes a critical limitation for complex transformation or geometry-intensive edits—like rotations or perspective shifts—where 2D-only reasoning leads to structural artifacts, unnatural deformations, and spatial misalignment. For example, plane-aware strategy decays displacement strength based solely on pixel distance, and ignores 3D geometry, resulting in perceptual distortions of human’s face as shown in Fig. 1(a, left).
@@
To improve the realism and semantic consistency of image editing, incorporating 3D geometric information is essential, as it offers richer structural cues beyond the 2D pixel plane. However, this introduces three key challenges. 1) How can geometry be integrated into pixel-level editing? While 3D cues (e.g., depth maps) are informative, they do not align directly with pixel-wise operations. For example, a drag defined in the image plane may become ambiguous in 3D due to perspective and depth variation—highlighting the need for a mechanism to incorporate geometry into 2D editing pipelines. 2) Is geometry alone sufficient for high-quality editing? Although helpful for preserving global structure, geometry alone may cause issues. Displacement fields based solely on 3D geometry (e.g., depth) often become discontinuous near object boundaries (see Sec. 3.1), disrupting the diffusion process and causing semantic artifacts as shown in Fig. 1(b). 3) How to reconcile guidance from multiple handle-target pairs? In real scenarios, users often specify multiple drag points. If their displacement fields overlap—especially with opposing directions—they can destructively interfere, even with distance-based weighting as in FastDrag [18]. This leads to displacement cancellation and failed edits as illustrated by Fig. 1(c). These challenges call for a unified framework that integrates 3D geometry and 2D cues while resolving conflicts for precise and coherent manipulation.
@@
Contributions. To tackle these challenges, we propose GeoDrag—a drag-based image editing framework that is both geometry-aware and plane-aware, ensuring coherent, high-fidelity, and fast one-step image manipulation. Built upon the latent consistency model (LCM) [20], GeoDrag predicts a dense displacement field directly in the noisy latent space at a specific diffusion timestep—circumventing iterative optimization and enabling fast and efficient editing. GeoDrag introduces three key innovations to resolve the above three challenges, respectively. 1) Geometry-aware field modeling: To resolve the mismatch between 3D geometry and 2D editing, GeoDrag introduces a novel influence function that modulates displacement strength based on 3D geometric relationships. By leveraging depth contrast, it ensures that nearby regions undergo stronger projective motion, while distant areas move more subtly—preserving 3D structure (see the well-edited human face in Fig. 1(a, right)). 2) Spatial plane modulation: Addressing the limitations of using 3D geometry alone, GeoDrag incorporates a spatial influence function based on 2D pixel plane. This improves local structure preservation and editing precision, especially in flat or geometry-ambiguous regions. 3) Conflict-Free Partitioning: To mitigate conflicts in multi-point editing, GeoDrag segments the editing mask into non-overlapping sub-regions, each associated with its nearest handle point. Independent displacement fields are computed per region, avoiding destructive interference and ensuring coherent multi-point manipulation. Together, these contributions allow GeoDrag to perform fast, high-quality, and semantically consistent edits in a single step—advancing the state of controllable, geometry-aware image manipulation.
@@ 2 Related Work
Text-Based Image Editing. Text-based image editing manipulates images via natural language prompts. Gal at al. [21] utilize textual inversion for personalized generation by embedding user-specific concepts. DiffusionCLIP [22] fine-tunes diffusion models with CLIP supervision, while Prompt-to-Prompt [4] achieves train-free editing by modifying cross-attention maps. Null-text inversion [2] optimizes unconditional text embeddings for faithful reconstruction of real images. Imagic [5] interpolates between text and image-specific embeddings but requires per-task tuning. Other approaches like CycleDiffusion [23] and DDPM inversion [9] explore latent spaces to support high-quality editing. InstructPix2Pix [24] trains on instruction-image pairs, enabling direct prompt-driven editing. RPG [6] introduces a multimodal LLM for reasoning and planning editing. Despite their flexibility, text-based methods often lack spatial precision and fine-grained control, limiting their applicability for detailed editing tasks.
Interactive Point-Based Image Editing. Point-based methods enable precise manipulation by directly dragging image elements. DragGAN [11] introduced this paradigm with GANs, later improved by diffusion-based approaches [25, 26, 20] like DragDiffusion [3], which integrates motion supervision and identity-preserving fine-tuning. Extensions enhance usability (EasyDrag [14]), stability (FreeDrag [13]), semantic control (DragNoise [12]), or robustness (GoodDrag [15], StableDrag [27]). To boost efficiency, FastDrag [18], RegionDrag [19], and SDEDrag [28] use lightweight latent manipulations, while DragonDiffusion [29], DiffEditor [30], and InstantDrag [17] introduce energy-based or real-time formulations. Despite progress, they remain limited to 2D pixel reasoning, restricting realism in geometry-sensitive edits. We address this with GeoDrag for controllable, structure-preserving, and efficient image editing. In parallel, FlowDrag [31] uses mesh reconstruction and iterative deformation to improve editing quality, but at a higher computational cost, that limits its responsiveness. By contrast, GeoDrag achieves geometry-consistent control while remaining responsive for interactive editing.
@@ 3 Methodology
We begin by formalizing the task of interest. Given an image and a set of point pairs , where each is a handle point and its corresponding target, the goal is to move each toward while preserving semantic consistency and visual realism. Recent efficient methods such as FastDrag [18] and RegionDrag [19] enable a fast one-step editing by computing a displacement field , which warps the latent at timestep via forward mapping:
@@
Existing methods [18, 19, 28] construct using only 2D pixel-plane heuristics, ignoring the underlying 3D structure. This estimated often leads to unrealistic deformations, semantic breaks, and perspective issues—especially in geometry-sensitive edits like rotations or viewpoint shifts (see Fig. 1 (a)). Additionally, multiple handle-target pairs can create overlapping and conflicting displacement fields, causing inconsistent guidance and editing failures (see Fig. 1 (c)).
@@ 3.2 Spatial Plane Modulation
Our fusion mechanism draws inspiration from elastic force propagation: deformation peaks at the force point and decays with distance. Mimicking this behavior, we define a plane-aware field that decays spatially from the handle point. This formulation allows localized and responsive editing, especially near fine structural features. While this idea is conceptually similar to the influence decay used in FastDrag [18], our method avoids plane geometric constructions like similar triangles and instead adopts a simpler, vectorized formulation that is more computationally efficient and easier to integrate.
@@ 4 Experiments
| DragDiffusion [3] | 34.57 | 0.181 | 0.170 | 0.160 | 0.871 | 1 min (LoRA) | 22.46 | 18.63 |
| FreeDrag [13] | 30.80 | 0.183 | 0.166 | 0.151 | 0.845 | 1 min (LoRA) | 42.90 | 18.90 |
| CLIPDrag [32] | 34.62 | 0.195 | 0.174 | 0.158 | 0.891 | 1 min (LoRA) | 38.21 | 22.72 |
| AdaptiveDrag [16] | 32.38 | 0.180 | 0.154 | 0.146 | 0.830 | 1 min (LoRA) | 46.30 | 7.71 |
| DragNoise [12] | 33.84 | 0.179 | 0.169 | 0.158 | 0.861 | 1 min (LoRA) | 21.12 | 18.36 |
| FastDrag [18] | 32.10 | 0.131 | 0.123 | 0.115 | 0.850 | ✗ | 3.23 | 5.85 |
@@ 4.1 Qualitative Evaluation
We conduct qualitative comparisons against existing state-of-the-art drag-based image editing methods, including DragDiffusion [3], DragNoise [12], FreeDrag [13], and FastDrag [18]. As shown in Fig. 5, GeoDrag achieves superior performance and high image quality. For instance, in the first row of Fig. 5, GeoDrag accurately drags the handle points toward the target points, preserving both structural integrity and semantic coherence, while other methods, such as FastDrag and FreeDrag, fail to maintain precise alignment. In multi-point editing (e.g., the second, fourth, and fifth rows of Fig. 5), GeoDrag successfully reshapes the wings, adjusts postures in alignment with the user-specified editing intention. In contrast, due to conflicts among multiple drag points, other methods struggle to generate coherent deformations. Observation from the last three rows of Fig. 5, GeoDrag produces 3D structure-coherent manipulation. These advantages are derived from the integration of geometric information, which provides a displacement field toward 3D structure alignment, enabling structurally consistent editing.
@@ 4.2 Quantitative evaluation
Here we conduct quantitative evaluations to validate GeoDrag where its modulation factors , , and are set to . Apart from [3, 12, 13, 18], CLIPDrag [32] and AdaptiveDrag [16] are included for comparison. DragBench dataset [3] is used as the benchmark, while Mean Distance (MD) [11] and Image Fidelity (IF) [5] are metrics for evaluating the editing precision and similarity between edited and original images. Dragging Accuracy Index (DAI) [15] measures consistency in the dragged region, with DAIr evaluating a radius- patch. We also report average editing time per point and peak GPU memory.
As reported in Table 1, GeoDrag achieves the best editing precision with the lowest MD and DAI, while maintaining competitive perceptual quality reflected by IF metric. Despite not requiring any preparation overhead, such as LoRA tuning, GeoDrag outperforms all baselines. GeoDrag edits each point in 3.95 seconds on average—faster than most diffusion-based methods—and consumes little GPU memory, making it suitable for responsive applications. More results on scalability w.r.t. the number of drag points are in Appendix F. We also conduct a user study using 10 randomly selected images, each edited by FreeDrag [13], FastDrag [18], and our GeoDrag. 60 Participants are asked to rank the edited results (1 for best, 3 for worst). As shown in Fig. 7, GeoDrag is superior to other methods.

### v2

@@ Abstract
Interactive point-based image editing serves as a controllable editor, enabling precise and flexible manipulation of image content. However, most drag-based methods operate primarily on the 2D pixel plane with limited use of 3D cues. As a result, they often produce imprecise and inconsistent edits, particularly in geometry-intensive scenarios such as rotations and perspective transformations. To address these limitations, we propose a novel geometry-guided drag-based image editing method—GeoDrag, which addresses three key challenges: 1) incorporating 3D geometric cues into pixel-level editing, 2) mitigating discontinuities caused by geometry-only guidance, and 3) resolving conflicts arising from multi-point dragging. Built upon a unified displacement field that jointly encodes 3D geometry and 2D spatial priors, GeoDrag enables coherent, high-fidelity, and structure-consistent editing in a single forward pass. In addition, a conflict-free partitioning strategy is introduced to isolate editing regions, effectively preventing interference and ensuring consistency. Extensive experiments across various editing scenarios validate the effectiveness of our method, showing superior precision, structural consistency, and reliable multi-point editability. Project page: https://xinyu-pu.github.io/projects/geodrag.
@@ 1 Introduction
Image editing (Cho et al., 2024; Mokady et al., 2022; Shi et al., 2024) has seen remarkable progress in recent years, largely driven by the emergence of powerful generative models (Hertz et al., 2022; Kawar et al., 2023; Mokady et al., 2022; Yang et al., 2024b; Zhou et al., 2025). Among these, text-guided image editing (Ruiz et al., 2023; Huberman-Spiegelglas et al., 2024; Zhou et al., 2025; Nguyen et al., 2024) has become a widely adopted paradigm, allowing users to modify images using natural language prompts. While expressive and flexible, this approach often falls short in providing fine-grained spatial control—especially when precise, localized, or geometry-sensitive edits are required (Pan et al., 2023; Shi et al., 2024; Liu et al., 2024; Ling et al., 2024; Hou et al., 2024; Zhang et al., 2025; Chen et al., 2024).
@@
To overcome these limitations, point-based image editing (Pan et al., 2023; Shi et al., 2024; Liu et al., 2024; Ling et al., 2024; Hou et al., 2024; Zhang et al., 2025; Chen et al., 2024) has emerged as a powerful alternative, gaining traction for its user-friendly, precise manipulation. By enabling users to specify handle-to-target point pairs, methods such as DragGAN (Pan et al., 2023) and its successors (Shi et al., 2024; Liu et al., 2024; Ling et al., 2024; Hou et al., 2024; Zhang et al., 2025; Chen et al., 2024; Shin et al., 2024) offer intuitive and precise image manipulation. They often adopt a two-step pipeline: (1) motion supervision, dragging the handle point toward the target, and (2) point tracking, monitoring the handle’s updated position during editing.
@@
Motivation. Despite their success, existing point-based methods often rely on iterative gradient-based optimization, which is computationally intensive and impractical for real-time use. To improve efficiency, recent approaches like FastDrag (Zhao et al., 2024) and RegionDrag (Lu et al., 2024) propose one-step editing via latent relocation based on dense displacement fields constructed over user-specified regions. However, these methods operate purely in the 2D pixel plane, and ignore the underlying 3D scene geometry. This becomes a critical limitation for complex transformation or geometry-intensive edits—like rotations or perspective shifts—where 2D-only reasoning leads to structural artifacts, unnatural deformations, and spatial misalignment. For example, plane-aware strategy decays displacement strength based solely on pixel distance, and ignores 3D geometry, resulting in perceptual distortions of human’s face as shown in Fig. 1(a, left).
@@
To improve the realism and semantic consistency of image editing, incorporating 3D geometric information is essential, as it offers richer structural cues beyond the 2D pixel plane. However, this introduces three key challenges. 1) How can geometry be integrated into pixel-level editing? While 3D cues (e.g., depth maps) are informative, they do not align directly with pixel-wise operations. For example, a drag defined in the image plane may become ambiguous in 3D due to perspective and depth variation—highlighting the need for a mechanism to incorporate geometry into 2D editing pipelines. 2) Is geometry alone sufficient for high-quality editing? Although helpful for preserving global structure, geometry alone may cause issues. Displacement fields based solely on 3D geometry (e.g., depth) often become discontinuous near object boundaries (see Sec. 3.1), disrupting the diffusion process and causing semantic artifacts as shown in Fig. 1(b). 3) How to reconcile guidance from multiple handle-target pairs? In real scenarios, users often specify multiple drag points. If their displacement fields overlap—especially with opposing directions—they can destructively interfere, even with distance-based weighting as in FastDrag (Zhao et al., 2024). This leads to displacement cancellation and failed edits as illustrated by Fig. 1(c). These challenges call for a unified framework that integrates 3D geometry and 2D cues while resolving conflicts for precise and coherent manipulation.
@@
Contributions. To tackle these challenges, we propose GeoDrag—a drag-based image editing framework that is both geometry-aware and plane-aware, ensuring coherent, high-fidelity, and fast one-step image manipulation. Built upon the latent consistency model (LCM) (Luo et al., 2023), GeoDrag predicts a dense displacement field directly in the noisy latent space at a specific diffusion timestep—circumventing iterative optimization and enabling fast and efficient editing. GeoDrag introduces three key innovations to resolve the above three challenges, respectively. 1) Geometry-aware field modeling: To resolve the mismatch between 3D geometry and 2D editing, GeoDrag introduces a novel influence function that modulates displacement strength based on 3D geometric relationships. By leveraging depth contrast, it ensures that nearby regions undergo stronger projective motion, while distant areas move more subtly—preserving 3D structure (see the well-edited human face in Fig. 1(a, right)). 2) Spatial plane modulation: Addressing the limitations of using 3D geometry alone, GeoDrag incorporates a spatial influence function based on 2D pixel plane. This improves local structure preservation and editing precision, especially in flat or geometry-ambiguous regions. 3) Conflict-Free Partitioning: To mitigate conflicts in multi-point editing, GeoDrag segments the editing mask into non-overlapping sub-regions, each associated with its nearest handle point. Independent displacement fields are computed per region, avoiding destructive interference and ensuring coherent multi-point manipulation. Together, these contributions allow GeoDrag to perform fast, high-quality, and semantically consistent edits in a single step—advancing the state of controllable, geometry-aware image manipulation.
@@ 2 Related Work
Text-Based Image Editing. Text-based image editing manipulates images via natural language prompts. Gal et al. (2023) utilize textual inversion for personalized generation by embedding user-specific concepts. DiffusionCLIP (Kim et al., 2022) fine-tunes diffusion models with CLIP supervision, while Prompt-to-Prompt (Hertz et al., 2022) achieves train-free editing by modifying cross-attention maps. Null-text inversion (Mokady et al., 2022) optimizes unconditional text embeddings for faithful reconstruction of real images. Imagic (Kawar et al., 2023) interpolates between text and image-specific embeddings but requires per-task tuning. Other approaches like CycleDiffusion (Wu and la Torre, 2023) and DDPM inversion (Huberman-Spiegelglas et al., 2024) explore latent spaces to support high-quality editing. InstructPix2Pix (Brooks et al., 2023) trains on instruction-image pairs, enabling direct prompt-driven editing. RPG (Yang et al., 2024b) introduces a multimodal LLM for reasoning and planning editing. Despite their flexibility, text-based methods often lack spatial precision and fine-grained control, limiting their applicability for detailed editing tasks.
Interactive Point-Based Image Editing. Point-based methods enable precise manipulation by directly dragging image elements. DragGAN (Pan et al., 2023) introduced this paradigm with GANs, later improved by diffusion-based approaches (Ho et al., 2020; Rombach et al., 2022; Luo et al., 2023) like DragDiffusion (Shi et al., 2024), which integrates motion supervision and identity-preserving fine-tuning. Extensions enhance usability (EasyDrag (Hou et al., 2024)), stability (FreeDrag (Ling et al., 2024)), semantic control (DragNoise (Liu et al., 2024)), or robustness (GoodDrag (Zhang et al., 2025), StableDrag (Cui et al., 2024)). To boost efficiency, FastDrag (Zhao et al., 2024), RegionDrag (Lu et al., 2024), and SDEDrag (Nie et al., 2024) use lightweight latent manipulations, while DragonDiffusion (Mou et al., 2024b), DiffEditor (Mou et al., 2024a), and InstantDrag (Shin et al., 2024) introduce energy-based or real-time formulations. Despite progress, they remain limited to 2D pixel reasoning, restricting realism in geometry-sensitive edits. We address this with GeoDrag for controllable, structure-preserving, and efficient image editing. In parallel, FlowDrag (Koo et al., 2025) uses mesh reconstruction and iterative deformation to improve editing quality, but at a higher computational cost, that limits its responsiveness. By contrast, GeoDrag achieves geometry-consistent control while remaining responsive for interactive editing.
@@ 3 Methodology
We begin by formalizing the task of interest. Given an image and a set of point pairs , where each is a handle point and its corresponding target, the goal is to move each toward while preserving semantic consistency and visual realism. Recent efficient methods such as FastDrag (Zhao et al., 2024) and RegionDrag (Lu et al., 2024) enable a fast one-step editing by computing a displacement field , which warps the latent at timestep via forward mapping:
@@
Existing methods (Zhao et al., 2024; Lu et al., 2024; Nie et al., 2024) construct using only 2D pixel-plane heuristics, ignoring the underlying 3D structure. This estimated often leads to unrealistic deformations, semantic breaks, and perspective issues—especially in geometry-sensitive edits like rotations or viewpoint shifts (see Fig. 1 (a)). Additionally, multiple handle-target pairs can create overlapping and conflicting displacement fields, causing inconsistent guidance and editing failures (see Fig. 1 (c)).
@@ 3.2 Spatial Plane Modulation
Our fusion mechanism draws inspiration from elastic force propagation: deformation peaks at the force point and decays with distance. Mimicking this behavior, we define a plane-aware field that decays spatially from the handle point. This formulation allows localized and responsive editing, especially near fine structural features. While this idea is conceptually similar to the influence decay used in FastDrag (Zhao et al., 2024), our method avoids plane geometric constructions like similar triangles and instead adopts a simpler, vectorized formulation that is more computationally efficient and easier to integrate.
@@ 4 Experiments
| DragDiffusion (Shi et al., 2024) | 34.57 | 0.181 | 0.170 | 0.160 | 0.871 | 1 min (LoRA) | 22.46 | 18.63 |
| FreeDrag (Ling et al., 2024) | 30.80 | 0.183 | 0.166 | 0.151 | 0.845 | 1 min (LoRA) | 42.90 | 18.90 |
| CLIPDrag (Jiang et al., 2024) | 34.62 | 0.195 | 0.174 | 0.158 | 0.891 | 1 min (LoRA) | 38.21 | 22.72 |
| AdaptiveDrag (Chen et al., 2024) | 32.38 | 0.180 | 0.154 | 0.146 | 0.830 | 1 min (LoRA) | 46.30 | 7.71 |
| DragNoise (Liu et al., 2024) | 33.84 | 0.179 | 0.169 | 0.158 | 0.861 | 1 min (LoRA) | 21.12 | 18.36 |
| FastDrag (Zhao et al., 2024) | 32.10 | 0.131 | 0.123 | 0.115 | 0.850 | ✗ | 3.23 | 5.85 |
@@ 4.1 Qualitative Evaluation
We conduct qualitative comparisons against existing state-of-the-art drag-based image editing methods, including DragDiffusion (Shi et al., 2024), DragNoise (Liu et al., 2024), FreeDrag (Ling et al., 2024), and FastDrag (Zhao et al., 2024). As shown in Fig. 5, GeoDrag achieves superior performance and high image quality. For instance, in the first row of Fig. 5, GeoDrag accurately drags the handle points toward the target points, preserving both structural integrity and semantic coherence, while other methods, such as FastDrag and FreeDrag, fail to maintain precise alignment. In multi-point editing (e.g., the second, fourth, and fifth rows of Fig. 5), GeoDrag successfully reshapes the wings, adjusts postures in alignment with the user-specified editing intention. In contrast, due to conflicts among multiple drag points, other methods struggle to generate coherent deformations. Observation from the last three rows of Fig. 5, GeoDrag produces 3D structure-coherent manipulation. These advantages are derived from the integration of geometric information, which provides a displacement field toward 3D structure alignment, enabling structurally consistent editing.
@@ 4.2 Quantitative evaluation
Here we conduct quantitative evaluations to validate GeoDrag where its modulation factors , , and are set to . Apart from (Shi et al., 2024; Liu et al., 2024; Ling et al., 2024; Zhao et al., 2024), CLIPDrag (Jiang et al., 2024) and AdaptiveDrag (Chen et al., 2024) are included for comparison. DragBench dataset (Shi et al., 2024) is used as the benchmark, while Mean Distance (MD) (Pan et al., 2023) and Image Fidelity (IF) (Kawar et al., 2023) are metrics for evaluating the editing precision and similarity between edited and original images. Dragging Accuracy Index (DAI) (Zhang et al., 2025) measures consistency in the dragged region, with DAIr evaluating a radius- patch. We also report average editing time per point and peak GPU memory.
As reported in Table 1, GeoDrag achieves the best editing precision with the lowest MD and DAI, while maintaining competitive perceptual quality reflected by IF metric. Despite not requiring any preparation overhead, such as LoRA tuning, GeoDrag outperforms all baselines. GeoDrag edits each point in 3.95 seconds on average—faster than most diffusion-based methods—and consumes little GPU memory, making it suitable for responsive applications. More results on scalability w.r.t. the number of drag points are in Appendix F. We also conduct a user study using 10 randomly selected images, each edited by FreeDrag (Ling et al., 2024), FastDrag (Zhao et al., 2024), and our GeoDrag. 60 Participants are asked to rank the edited results (1 for best, 3 for worst). As shown in Fig. 7, GeoDrag is superior to other methods.